# Deanonymizing Ethereum Validators: The P2P Network Has a Privacy Issue

### v1

@@ 1 Introduction
In this work, we demonstrate the impact of these scaling solutions on the privacy and security of the Ethereum P2P network and blockchain. Specifically, we outline how to deanonymize validators in the P2P network by mapping a validator’s identifier to the IP address of the associated consensus layer node. Our technique relies solely on observing attestation (i.e., consensus layer vote) messages received from peers (i.e., nodes with established TCP connections). By analyzing messages from a peer $p$, we can infer whether a validator $v$ is associated with that peer.
@@
The Ethereum P2P network’s privacy issue poses a major security risk, allowing attackers to identify nodes associated with validators set to create new blocks. This could lead to (D)DOS attacks, halting chain progress, or more targeted attacks on nodes associated with validators handling high-MEV blocks, letting a subsequent malicious validator scoop these profits. We hope our work highlights this lack of privacy and informs future privacy-enhancing solutions.
@@
  • We propose a simple and low-cost technique for a node in the network to deanonymize its peers, i.e., infer which validators are associated with them.
@@
  • From our measurements, we draw novel insights about the P2P network. In particular, we comment on the distribution of validators across peers in the P2P network (e.g., we locate more than 19,000 validators on one peer) and illustrate the distribution of validators across the globe and organizations (i.e., cloud service and internet service providers).
  • Finally, we outline the implication of the lack of anonymity in Ethereum’s P2P network (e.g., liveness and safety concerns) and discuss possible mitigations.
Responsible Disclosure We submitted a bug report to the Ethereum Foundation, disclosing the vulnerabilities presented in this paper. The Ethereum Foundation awarded us a bug bounty, acknowledging the impact of our results. We discuss further ethical considerations in Appendix A.
@@ 2.1 The Ethereum Blockchain
The Ethereum blockchain operates as a Proof-of-Stake blockchain. More precisely, individuals can pay 32 ETH to become a validator, i.e., a participant in the consensus algorithm. The 32 ETH represent a validator’s stake and can be confiscated in case of inactivity or malicious behavior. An individual can either run their own validator(s), i.e., solo/home stake or stake through a staking pool. Staking pools allow participants to pool resources to partake in the consensus algorithm with arbitrarily sized amounts of stake.
The consensus algorithm is divided into epochs, with each epoch lasting 32 slots [32]. A slot corresponds to one new block appended to the blockchain. Among other validity rules, each block must be created (or proposed) by the corresponding validator assigned to it and have sufficient attestations of its correctness from the group (or commitee) of validators assigned to it. We give an overview of relevant details below.
Validators. Three main duties fall on validators [30]. First, during each slot one validator is selected to propose a new block. Secondly, during each epoch, each validator is assigned to a committee in one slot to make an attestation. Finally, with some probability, a validator is chosen as an aggregator for a slot they are attesting, i.e., they are tasked with collecting attestations and publishing an aggregate. Validators are identified by an ID, linked to a private/public key pair. The logical entity described is hosted on a validator client, a separate entity with access to one or multiple validator private keys.
Nodes. A validator client interacts with the rest of the network through consensus nodes, also referred to as node in the following. A node manages its own identity, consisting of a public/private key pair, linked to an IP address and port number. This information is shared in the shape of Ethereum Node Records (ENR), which are propagated across the network. Importantly, a node can host any number of validators, as these two entities are kept fully separate, in part for security reasons (see Section 9). Moreover, a node might not host any validators, e.g., researchers who simply want to read the blockchain state.
@@ 2.2 Scaling Attestation Dispersal in Ethereum
The Ethereum P2P network protocol facilitates the exchange of messages – most importantly blocks, attestations, and attestation aggregations—among peers, which is required for the blockchain to progress.
@@
Gossip Protocol. A further network optimization to reduce message complexity is that messages are shared using a probabilistic broadcast implementation called GossipSub [65]. When a validator signs an attestation, the consensus node to which it is connected publishes the attestation to the corresponding subnet by sending it to a subset of peers that are part of this subnet, these are called the node’s fanout for this subnet. Note that the validator’s consensus client itself does not need to be subscribed to this subnet, as committee assignments for validators change every epoch.
@@
To ensure stability within a subnet, each node statically subscribes to two topics by default, performing backbone duties, which can be considered static subscriptions. Additionally, if nodes need to receive messages from non-subscribed subnets (e.g., for an upcoming aggregation duty), they can request a dynamic subscription. Within a subscribed subnet, nodes choose a subset of peers that are also in the subnet to share messages with. The choice of this subset is based on peer-performance***Performance is based on a peer-scoring function which takes into account several factors including speed and quantity of information forwarded [66]. Specifics on the behavior of this score function is out of the scope.. Nodes forward all messages they hear about within a subnet to these best-performing peers. These connections make up the subgraphs of Figure 1.
@@ 3 Deanonymization Methodology
Our goal is to locate all validators running on a peer we are connected to, and thereby link validator addresses to their IP address. Our method is simple in that it solely focuses on attestation messages and their dissemination subnets. We show these are sufficient for our deanonymization attack.
@@ 3.1 Ideal Approach
Given the background on Ethereum node behavior, we can describe how an ideal peer (a peer who gives us perfect information) would behave. We will guide the description with an example of a real peer we connected to in our experiments (see Figure 2). Let us assume we are connected to a peer running $V$ validators who is a backbone in***We use the terms subscribed to and backbone in interchangeably. two subnets. The peer’s validators will attest $V$ times per epoch. Let us assume we receive perfect information from this peer, meaning we are in their fanout for all subnets and they forward all attestations they hear about in their two backbones to us. In each epoch, we will receive $V$ attestations from our peer for their validators, and $N\cdot\frac{2}{64}$ for all other $N$ validators.
Figure 2: Attestations received from a peer over a four hour window, where the vertical axis corresponds to the 64 subnets. We identified four validators hosted on the machine, represented in red, blue, yellow, and green. The attestations by the remaining validators are grouped together in pink. Note that we have anonymized the full IDs of the identified validators, but emphasize that the IDs of these four validators are sequential. For the identified validators, we receive attestations from a wide variety of subnets. In contrast, the attestations from the remaining validators primarily come from the two subnets the peer predominantly serves (i.e., subnets 12 and 13), evidenced by the repeated flow of attestations for these subnets (see the long pink strip), as well as from dynamic subscriptions (see the smaller pink horizontal strips).
Observation 1. An ideal peer will only send us an attestation in a subnet they are not a backbone of if they are the signer of the attestation, and we are in their fanout for the corresponding subnet of the attestation.
@@
Imperfect Information in Practice. To showcase this, we turn to the example peer in Figure 2. On this peer, we will identify four validators associated with the peer;***Using out heuristics introduced in the following subsection. their respective attestations are highlighted in red, blue, yellow, and green, while the remaining attestations are shown in pink. Notice that the attestations from these four validators, who have consecutive identifiers, appear equally distributed across subnets. In contrast, the vast majority of attestations come from the two subnets where the peer acts as a backbone (subnets 12 and 13 for the sample peer in Figure 2). Thus, we can locate validators on our peers by observing how the attestations belonging to a validator, which we receive from the peer, are distributed across subnets. Additionally, and this is where the imperfect information comes into play, the validators hosted on the peer are occasionally tasked with being aggregators in a subnet approximately every 30 epochs per validator. During these times, they temporarily become a backbone (see the smaller pink horizontal strips) for these subnets and receive attestations from multiple validators belonging to the subnet.
In practice, five main factors impact the level of information we receive from a peer. First, nodes may run at non-default parameters, such as in more than two backbones, increasing the number of messages we receive from the peer. Second, we may not receive all attestations that a peer generates due to network reasons (e.g., disconnections or dropped network packets), or dynamic membership in a peer’s fanouts which is out of our control (a peer chooses whether to include us in a fanout or not). Third, due to dynamic subscriptions or delayed information, we may receive some backbone attestations from a peer and not label them as such (see Figuer 2). Fourth, a validator client might use multiple nodes to propagate their messages or use different nodes for separate tasks [5]. Finally, fifth, we only get propagated attestations from peers if they give us a good peer score, thus we have to participate in attestation propagation. Us announcing a subset of attestations to our peers first in each backbone results in fewer attestations from each peer. This should be rare if they are the original signer of the attestations, and thus we make the following assumption to guide the development of our heuristics:
Assumption 1. A peer will be the first to tell us about their attestation most of the time.
@@ 3.2 Heuristic Approach
To handle imperfect information, we develop a heuristics deanonymization approach to filter peers and narrow down on validator IDs as follows: we consider a validator $v$ to be hosted***As pointed out in Section 3.1 a validator client might use multiple nodes to propagate their messages. Thus, to put it differently, the node is associated with the validator and we will use the terms interchangeably throughout. We also observe such behavior, i.e., validators associated with multiple nodes, as we will outline in Section 6. However, this behavior is rare and likely only used by node operators of large staking pools given that running a consensus node is expensive. Further, our methodology in theory allows us to locate all consensus nodes associated with a validator. on a peer $p$, if all of the following conditions hold:
@@
    The number of attestations we receive for the validator $v$ from the peer $p$ exceeds the mean number of attestations per validator from peer $p$ by two standard deviations.
Condition 1 ensures that the proportion of attestations we receive from $p$’s fanouts for a validator $v$ is at least 90% the expected value. We choose this number (the 90%) conservatively to analyze only those peers from which we receive most attestations across all subnets***In robustness-testing we find that if we lower this to 70% we find an additional 1K validators overall, accounting for less than 0.4% additional validators. We leave more fine-grained approaches for future work.. Condition 2 excludes any validators that are in all backbones as we would get no non-backbone attestations. Due to our ephemeral membership in a peer’s fanouts and the backbone peers they broadcast to, condition 3 ensures that we receive sufficient data from our peer throughout the connection window***We use a conservative bound of at least one-tenth of messages across all subnets, though consistent attestations in just one fanout may already be enough to link validators.. Lastly, condition 4 increases our confidence that the peer is not participating in rare, non-default behavior such as propagating in all subnets without advertising their subscriptions.
@@ 4 Data Collection
We now detail how we collect data to both run our heuristic deanonymization approach and validate its results.

### v2

@@ 1 Introduction
Our work demonstrates the impact of these scaling solutions on the privacy and security of the Ethereum P2P network and blockchain. We outline how to deanonymize validators in the P2P network by mapping a validator’s identifier to the IP address of the machine it is hosted on. Our technique relies solely on observing attestation (i.e., consensus layer vote) messages received from peers (i.e., nodes with established TCP connections). By analyzing messages from a peer $p$, we can infer whether a validator $v$ is hosted on this peer.
@@
The Ethereum P2P network’s privacy issue poses a major security risk, allowing attackers to identify nodes associated with validators set to create new blocks. This could lead to (D)DoS attacks, halting chain progress, or more targeted attacks on nodes associated with validators handling high value blocks, letting a subsequent malicious validator scoop these profits. We hope our work highlights this lack of privacy and informs future privacy-enhancing solutions.
@@
  • We propose a simple and low-cost technique for a node in the network to deanonymize its peers, i.e., infer which validators they host.
@@
  • We outline the implication of the lack of anonymity in Ethereum’s P2P network (e.g., fairness, liveness and safety concerns) and discuss possible mitigations.
  • Finally, we expose novel security risks in the P2P network, highlighting how validators concentrate on certain peers (e.g., we locate over 19,000 validators on a single peer) and how they are spread globally and across organizations (i.e., cloud service and internet service providers). We also discover that operators for different staking pools run multiple pools’ validators on the same machine, creating undesirable dependencies.
Responsible Disclosure We submitted a bug report to the Ethereum Foundation, disclosing the vulnerabilities presented in this paper. The Ethereum Foundation awarded us a bug bounty, acknowledging the impact of our results. We discuss further ethical considerations in Section 12.
@@ 2.1 The Ethereum Blockchain
The Ethereum blockchain operates as a Proof-of-Stake blockchain, and consists of the consensus layer (also called the Beacon chain) and the execution layer. Individuals can pay 32 ETH to become a validator, i.e., a participant in the consensus algorithm. The 32 ETH represent a validator’s stake and can be confiscated in case of inactivity or malicious behavior. An individual can either run their own validator(s), i.e., solo/home stake or stake through a staking pool. Staking pools allow participants to pool resources to partake in the consensus algorithm with arbitrarily sized amounts of stake. [^1] Note that the staking pool is responsible for fully operating the validator(s), unlike in Proof-of-Work mining pools.
The consensus algorithm is divided into epochs, with each epoch lasting 32 slots [28]. A slot lasts 12 seconds, resulting in an epoch duration of 6 minutes and 24 seconds. A slot corresponds to one new block appended to the blockchain. Among other validity rules, each block must be created (or proposed) by the corresponding validator assigned to it and have sufficient attestations of its correctness from the group (or commitee) of validators assigned to it. We give an overview of relevant details below.
Validators. Three main duties fall on validators [27]. First, during each slot one validator is selected to propose a new block. Second, during each epoch, each validator is assigned to a committee in one slot to make an attestation. Finally, with some probability, a validator is chosen as an aggregator for a slot they are attesting, i.e., they are tasked with collecting attestations and publishing an aggregate. Validators are identified by an ID, linked to a private/public key pair. The logical entity described is hosted on a validator client, a separate entity with access to one or multiple validator private keys.
Nodes. A validator client interacts with the rest of the network through a consensus node, also referred to as node throughout. A node manages its own identity, consisting of a public/private key pair, linked to an IP address and port number. This information is shared through Ethereum Node Records (ENR), which are propagated across the network. Importantly, a node can host any number of validators, as these two entities are kept separate, in part for security reasons (see Section 9). Moreover, a node might not host any validators, e.g., researchers who simply want to read the blockchain state.
@@ 2.2 Scaling Attestation Dispersal in Ethereum
The Ethereum P2P network protocol facilitates the exchange of messages – most importantly blocks, attestations, and attestation aggregations – among peers, which is required for the blockchain to progress.
@@
Gossip Protocol. A further network optimization to reduce message complexity is that messages are shared using a probabilistic broadcast implementation called GossipSub [76]. When a validator signs an attestation, the consensus node to which it is connected publishes the attestation to the corresponding subnet by sending it to a subset of peers that are part of this subnet, these are called the node’s fanout for this subnet. Note that the sending consensus node itself does not need to be subscribed to this subnet, as committee assignments for validators change every epoch.
@@
To ensure stability within a subnet, each node statically subscribes to two topics by default, performing backbone duties, which can be considered static subscriptions. Additionally, if nodes need to receive messages from non-subscribed subnets (e.g., for an upcoming aggregation duty), they can request a dynamic subscription. Within a subscribed subnet, nodes choose a subset of peers that are also in the subnet to share messages with. The choice of this subset is based on peer-performance. [^2] Performance is based on a peer-scoring function which takes into account several factors including speed and quantity of information forwarded [77]. Specifics on the behavior of this score function is out of scope. Nodes forward all messages they hear about within a subnet to these best-performing peers. These connections make up the subgraphs of Figure 1.
3 Threat Model
In this work, we consider the P2P network of the Ethereum consensus layer, where consensus messages are broadcast. This system’s main goal is reliable, efficient and low latency message delivery among a dynamic set of nodes hosting validators. As a secondary goal, source anonymity is also desired [68]. Below we detail two main threats in the absence of source anonymity.
We consider an attacker aiming to create a wide-scale deanonymization, namely map validators to IP addresses of their hosting nodes. Importantly, since all validators have equal stake, no particular validator must be targeted.
We assume a weak, passive attacker, with access to a few nodes (without validators), limited time and hardware budget. The attacker only has access to exchanged network-level messages. All other data sources and attack vectors (off-chain data, clustering, etc.) are used for the verification of results.
Deanonymization is deemed successful if the probability that a given validator resides at a particular IP address is significantly higher than that expected from random assignment. The feasibility of deanonymizing a large portion of Ethereum’s validators within the P2P network presents significant security concerns, enabling various types of attacks. These threats include those detailed in the following, as well as others, such as undermining assumptions critical for Danksharding [63] (a new Ethereum upgrade) and compromising censorship resistance. Notably, even imperfect knowledge in these scenarios can inflict substantial harm on the network.
3.1 Taking-Out Preceding Block Proposers
Ethereum block rewards comprise both consensus layer rewards and execution layer rewards. The consensus layer rewards are constant and independent of the transactions included in a block. In contrast, execution layer rewards consist of transaction fees paid for (preferential) block inclusion. On average, execution layer rewards are over three times higher than consensus layer rewards and can surpass them by several orders of magnitude due to the presence of MEV (Maximal Extractable Value) in the Ethereum blockchain [41]. MEV is any value that can be extracted by the block builder by including, excluding, and reordering transactions in a block.
This disparity poses a risk to the consensus layer, as previous works [19, 67] have shown. Specifically, the consensus layer is vulnerable to time-bandit attacks, where it can be rational for a block proposer to fork the blockchain and extract the MEV from earlier blocks for personal gain. The traditional time-bandit attack involves forking out the block created by the previous proposer, which is more challenging in Ethereum’s PoS system than in PoW [48].
Critically, attackers can exploit the lack of privacy in the P2P network to enhance time-bandit attacks and bypass the need to fork out the previous proposer’s block.
An attacker, as the proposer of slot $n+1$, can prevent the proposer of slot $n$ from submitting a block proposal [17]. By doing so, the attacker could claim higher execution layer rewards, incorporating transactions from the previous slot and any new transactions arriving in the interim.
The attacker, as the proposer of slot $n+1$, knows the identity of their victim (the proposer of slot $n$) in the consensus layer well in advance, as proposers are assigned at least one epoch (approximately six minutes) beforehand. If the attacker can deanonymize the victim in the P2P network, they might temporarily sever the victim’s connection to the network through a (D)DoS attack or BGP hijacking. This disruption only needs to last four seconds, as block proposers have a four-second window to submit their proposal.
The success of this attack hinges on the attacker’s ability to both identify the victim in the P2P network and disconnect them. While some victims may have implemented mitigations (as outlined in Section 9) that prevent the attack, an attacker can simply wait for another opportunity when they are selected as the block proposer and attempt the attack again.
3.2 Breaking Liveness and Safety
Even more concerning is the possibility of an attacker escalating the previous attack to compromise the liveness or safety of Ethereum. To disrupt liveness, an attacker could repeatedly sever the connection between the upcoming block proposer and the network. If every such attack succeeded, the blockchain would come to a standstill. However, achieving consistent success is improbable due to the mitigation measures discussed in Section 9. Large staking pools, such as Lido, are particularly likely to implement robust defenses. Nevertheless, even if an attacker could disrupt just one in every ten proposers, the consequences for the blockchain would be significant, leading to delays and instability.
Rather than preventing a proposer from submitting a block entirely, the attacker might instead seek to disrupt block propagation to over one-third of the network. This could be achieved by launching a sustained DoS attack on a significant portion of the network or leveraging BGP hijacking. Such an attack would halt liveness, as the finality gadget – responsible for confirming blocks – would fail to gather the quorum required to finalize blocks. Again, the success of this strategy hinges on the attacker’s ability to prevent block proposals from reaching at least one-third of validators.
Alarmingly, an attacker could also threaten the safety of the blockchain. Many Ethereum light clients optimistically treat the chain’s head as finalized. By breaking synchrony assumptions through targeted DoS attacks or BGP hijacking, the attacker could exploit this optimistic behavior to introduce safety violations [20]. For example, conflicting blocks might appear finalized to different segments of the network, resulting in inconsistencies that undermine the integrity of the chain.
Although such attacks are complex and resource-intensive, they are made more feasible by the deanonymization techniques discussed in the following.
@@ 4 Deanonymization Methodology
To ensure that their attack is as successful as possible, an attacker would aim to locate all validators running on a peer they are connected to, and thereby link validator IDs to their IP address. We show that focusing on attestation messages and their dissemination suffices for our deanonymization attack.
@@ 4.1 Ideal Approach
Given the background on Ethereum node behavior, we can describe how an ideal peer (a peer who gives us perfect information) would behave. We will guide the description with an example of a real peer we connected to in our experiments (see Figure 2). Let us assume we are connected to a peer running $V$ validators who is a backbone in [^3] We use the terms subscribed to and backbone in interchangeably. two subnets. The peer’s validators will thus attest $V$ times per epoch. Let us assume we receive perfect information from this peer, meaning we are in their fanout for all subnets and they forward all attestations they hear about in their two backbones to us. In each epoch, we will receive $V$ attestations from our peer for their validators, and $N\cdot\frac{2}{64}$ for all other $N$ validators.
Figure 2: Attestations received from a peer over a four hour window, where the vertical axis corresponds to the 64 subnets. We identified four validators hosted on the machine, represented in red, blue, yellow, and green. The attestations by the remaining validators are shown in pink. Note that we have anonymized the full IDs of the identified validators, but emphasize that the IDs of these four validators are sequential. For the identified validators, we receive attestations from a wide variety of subnets. In contrast, the attestations from the remaining validators primarily come from the two subnets the peer predominantly serves (i.e., subnets 12 and 13), evidenced by the repeated flow of attestations for these subnets (see the long pink strip), as well as from short dynamic subscriptions (see the smaller pink horizontal strips).
Observation. An ideal peer will only send us an attestation in a subnet they are not a backbone of if they are the signer of the attestation, and we are in their fanout for the corresponding subnet of the attestation.
@@
Imperfect Information in Practice. To showcase this, we turn to the example peer in Figure 2. On this peer, we will identify four validators associated with the peer; [^4] Using our heuristics introduced in the following subsection. their respective attestations are highlighted in red, blue, yellow, and green, while the remaining attestations are shown in pink. Notice that the attestations from these four validators, who have consecutive identifiers, appear equally distributed across subnets. In contrast, the vast majority of attestations come from the two subnets where the peer acts as a backbone (subnets 12 and 13 for the sample peer in Figure 2). Thus, we can locate validators on our peers by observing how the attestations belonging to a validator, which we receive from the peer, are distributed across subnets. Additionally, and this is where the imperfect information comes into play, the validators hosted on the peer are occasionally tasked with being aggregators in a subnet approximately every 30 epochs per validator. During these times, they temporarily become a backbone (see the smaller pink horizontal strips) for these subnets and receive attestations from multiple validators belonging to the subnet.
In practice, five main factors impact the level of information we receive from a peer. First, nodes may run at non-default parameters, such as in more than two backbones, increasing the number of messages we receive from the peer. Second, we may not receive all attestations that a peer generates due to network reasons (e.g., disconnections or dropped network packets), or dynamic membership in a peer’s fanouts which is out of our control (a peer chooses whether to include us in a fanout or not). Third, due to dynamic subscriptions or delayed information, we may receive some backbone attestations from a peer and not label them as such (see Figure 2). Fourth, a validator client might use multiple nodes to propagate their messages or use different nodes for separate tasks [57]. Finally, fifth, we only get propagated attestations from peers if they give us a good peer score, thus we have to participate in attestation propagation. Us announcing a subset of attestations to our peers first in each backbone results in fewer attestations from each peer. This should be rare if they are the original signer of the attestations, and thus we make the following assumption to guide the development of our heuristics.
Assumption. A peer will be the first to tell us about their attestation most of the time.
@@ 4.2 Heuristic Approach
To handle imperfect information, we develop a heuristics deanonymization approach to filter peers and narrow down on validator IDs as follows: we consider a validator $v$ to be hosted [^5] As pointed out in Section 4.1 a validator client might use multiple nodes to propagate their messages. Thus, to put it differently, the node is associated with the validator and we will use the terms interchangeably throughout. We also observe such behavior, i.e., validators associated with multiple nodes, as we will outline in Section 7. However, this behavior is rare and likely only used by node operators of large staking pools given that running a consensus node is expensive. Further, our methodology in theory allows us to locate all consensus nodes associated with a validator. on a peer $p$, if all of the following conditions hold:
@@
    The number of attestations we receive for the validator $v$ from the peer $p$ exceeds the mean number of attestations per validator from peer $p$ by one standard deviation.
By participating in all subnets w.h.p. we are added to several fanouts of our peer.  [^6] We discuss this likelihood further in Appendix A.
Condition 1 ensures that the proportion of attestations we receive from $p$’s fanouts for a validator $v$ is at least 90% the expected value. We choose 90% conservatively to analyze only those peers from which we receive most attestations across all subnets. [^7] In robustness-testing we find that if we lower this to 30% we find an additional 10K validators overall, accounting for less than 4% new validators.
Condition 2 excludes any validators that are in all backbones as we would get no non-backbone attestations. Condition 3 ensures that we analyze the peers for which we are in several fanouts, removing those that would require a more careful analysis (we leave these peers unlabeled). [^8] We use a conservative bound of at least one-tenth of messages across all subnets, though consistent attestations in just one fanout may already be enough to link validators. Lastly, condition 4 disregards peers that are participating in rare, non-default behavior such as broadcasting attestations in all subnets without advertising their subscriptions.
The above conditions represent conservative heuristics such that we analyze only peers for which our connections were stable enough, for long enough, to ensure we have sufficient data. In Appendix A, we provide a detailed explanation of all heuristics and robustness testing for each parameter choice.
@@ 5 Data Collection
We provide an in-detail explanation of how we collect data to both run our heuristic deanonymization approach and validate its results.